OpenStudy (anonymous):

Let y = (6 - 2 x)^4. Find the first and second derivatives of y. dy/dx= A ( 6 - 2x )B d2y /dx2 = C ( 6 - 2x )D

sam (.sam.):

A=4 B=-2 C=-24 D=-2

OpenStudy (anonymous):

i put that in for a= 4 and said it was wrong

sam (.sam.):

y'= 4(6-2x)^3 (-2) y'= -8(6-2x)^3 y''= -24(6-2x)^2 (-2) y''=48(6-2x)^2

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

use chain rule (6 - 2 x)^4 g(x) = x^4 g'(x) = 4x^(3) f(x) = 6-2x f'(x) = -2 y' = (4(6-2x)^(3)) (-2) =-8(6-2x)^(3) use chain rule again g(x) = -8x^3 g'(x) = -8(3)x^(2) f(x) = 6-2x f'(x) = -2 y'' = (-8(3)(6-2x)^(2))(-2) y'' = 48(6-2x)^(2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Derivatives are really easy to understand especially chain rule if you split them up like I have

OpenStudy (anonymous):

A = 4, B = -2, C = -24, D = -2 y = (6 - 2 x)^4. y' = 4 ((6 - 2 x)^3 ( -2) = -8 ((6 - 2 x)^3 y" = - 24 (6 - 2 x)^2 ( -2) = 48 (6 - 2 x)

OpenStudy (anonymous):

I means = 48 (6 - 2 x)^2